OpenStudy (anonymous):

Given that (5,4) is on the graph of f(x), find the corresponding point for the function f(-1/3x).

OpenStudy (jtvatsim):

This question is phrased a little strangely which makes it hard to understand exactly what they want. But consider this, 1. We know that (5,4) is on the graph of f(x). That is, "When x = 5, then y = 4." 2. This means that f(5) = 4. In English, "When x = 5, then y = 4." (same thing) 3. We are to find the "corresponding point" for f(-1/3x). This sounds weird. 4. But, all we know is that f(5) = 4. This is the only information we have. 5. But look! I can be clever and do this... 6. Take f(-1/3 x) and plug in x = -15. Then, 7. f(-1/3 * -15) = f(5) = 4. 8. So the corresponding point must be x = -15 and y = 4. That is, (-15, 4).

OpenStudy (anonymous):

@jtvatsim You're a life saver. I'm still slightly confused on this but you explained this very well. Could you help me solve another one?

OpenStudy (jtvatsim):

Glad to help in any little way. If you had any questions on this one, just let me know. It had me stumped for awhile myself... :) Feel free to post a new question on the site so that everyone can see it.